OpenStudy (anonymous):

HELP I WILL GIVE MEDALS !! What is the location of the positive focus point of this ellipse? x/225+y/216= 1

OpenStudy (anonymous):

\[\frac{ x^2 }{ 15^2 }+\frac{ y^2 }{ \left( \sqrt{216} \right)^2 }=1\] compare with \[\frac{ x^2 }{ a^2 }+\frac{ y^2 }{ b^2 }=1\] \[b^2=a^2\left( 1-e^2 \right),216=225\left( 1-e^2 \right),-e^2=\frac{ 216-225 }{ 225 },e=\pm \frac{ 1 }{ 5 }\] positive focus is (ae,0)=? (take only positive sign)

OpenStudy (anonymous):

216 isn't squared.. Would it still be the same? The answer has to be in numbers like (5,6)

OpenStudy (anonymous):

\[\frac{ x^{2} }{ 25 }+\frac{ y ^{2} }{ 16 }\]

OpenStudy (anonymous):

I think it put the question wrong.. It should have looked like that @surjithayer